The length of a rectangle is six inches less than three times the width.The perimeter of the rectangle is 36 inches.What are the width and length of the rectangle?





PLEASE SHOW YOUR WORK.tHIS IS A TEST IN MY GEOMETRY CLASS THAT HE LET US TAKE HOME.

Answer by checkley77(12844)

L=3W-6
2L+2W=P
2(3W-6)+2W=36
6W-12+2W=36
8W=36+12
6W=48
W=48/6
W=8 FOR THE WIDTH.
L=2*8-6
L=16-6
L=10
PROOF;
2*10=2*8=36
20+16=36
36=36